# The Loser’s Curse and the Critical Role of the Utility Function

### v3

@@ 4 Research hypothesis
Massey and Thaler, (2013) pioneered the traditional analytical approach to constructing draft position value curves. Their study produced two key findings. First, they discovered the loser’s curse––last years worst teams are “cursed” with the ostensible gifts of the top picks, which cost a lot (in compensation) without delivering enough on-field value to compensate. Second, they demonstrated that the draft pick trade market is highly inefficient––if one adopts the expected surplus value utility function, teams significantly overvalue top draft picks, consistently overpaying to trade up in the draft without trading down often enough.
Despite the widespread dissemination of these findings, general managers’ draft behavior has remained largely unchanged. When applied to recent data (2013––2023), the loser’s curse persists and the fitted trade market curve remains similar (see Section 3 and Avery et al., ming, forthcoming). If general managers were optimizing for expected performance value or expected surplus value, their continued reluctance to trade down and their willingness to pay a premium for higher picks would suggest systematic misjudgment. One possible explanation is that general managers have simply failed to learn from these analyses. An alternative explanation is that general managers are instead operating under a fundamentally different utility function. If maximizing expected surplus value does not align with their actual objectives, then their observed behavior may be rational, even if it appears suboptimal under Massey and Thaler’s framework. This paper challenges their core assumption, that general managers should adopt expected surplus value as their utility function in valuing draft picks. This assumption is neither explicitly justified in their analysis nor necessarily optimal in the context of NFL decision making.
@@ 5.1 Position-agnostic conditional density
In Figure 4 we visualize the empirical conditional density of performance $Y$ given draft position $x$. The empirical densities are extremely wobbly and noisy due to limited data for each individual draft position. But there is a general trend: we see a spike near zero that grows upwards as draft position increases and a right tail that shifts leftwards and diminishes as draft position increases. This motivates our “Bayesian spike plus Beta regression” model of ${\mathbb{P}}(Y|x)$.111111 We initially tried a lone Beta regression model without the spike near zero, which less successfully fit the right tail.
@@
We model the spike near zero, the bust spike, using logistic regression. We let a bust be a player whose performance outcome $Y$ is smaller than some cutoff ${y_{\text{bust}}}$. We found that ${y_{\text{bust}}}=0.5\%$ of the cap works well. We model the bust probability ${\mathrm{bp}}(x):={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$ by
@@
where $\widetilde{x}$ is a B-spline basis.121212 The ${\mathsf{R}}$ code we used to generate the spline basis is splines::bs($x$, degree=3, df=4, intercept=T). We model the conditional precision by
@@
To capture uncertainty in the parameters, we use a Bayesian model, placing priors on the parameters. We put a weakly informative Normal prior with mean zero and standard deviation ten on each of the $\alpha$, $\beta$, and $\gamma$ parameters. Finally, we assume that $Y$ is uniformly distributed within the bust spike in a way that doesn’t vary with $x$: $Y|x,Y\leq{y_{\text{bust}}}\sim{\mathrm{Unif}}[0,{y_{\text{bust}}}]$.131313 We assess the reasonableness of this assumption later in this section.
@@
Figure 5: Left: our estimate of the conditional mean $x\mapsto\mu(x)={\mathbb{E}}[Y|x,Y>{y_{\text{bust}}}]$ given not a bust. Middle: our estimate of the conditional standard deviation $x\mapsto{\mathrm{sd}}(x)={\mathrm{sd}}(Y|x,Y>{y_{\text{bust}}})$ given not a bust. Right: our estimate of bust probability $x\mapsto{\mathrm{bp}}(x)={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$. The solid black lines are the posterior means, the cyan shaded regions are the $95\%$ credible intervals, and the gray dots are the empirical mean values for each draft position $x$.
Figure 6: The distribution of performance $Y$ ($x$-axis) given draft position $x$ (facet). The solid black lines are the fitted conditional densities ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})$ (the posterior mean curves) and the cyan shaded regions are the $95\%$ credible intervals, which overlay the empirical distributions (histogDolphins and smoothed densities) in gray.
@@ 6 Results
Given our estimated conditional density of performance given draft position and player position, in this section we construct draft position value curves from alternative objective functions. We move beyond expected performance value and expected surplus value to objective functions that emphasize elite, right-tail outcomes.
@@ 6.1 Right tail probability
Further, we find that general managers trade draft picks as if they value a draft position by the probability it produces a player whose second contract exceeds $19.7\%$ of the salary cap. Specifically, the mean absolute error between the fitted trade market curve and a right tail probability curve with eliteness threshold $r$ is minimized by $r=19.7\%$ (note that Deshaun Watson has $r=19.7\%$). This is a high bar of eliteness: among quarterbacks in our dataset who secured multi-year second contracts as regular starters, just $12.3\%$ of them ($8/65$) met this criterion. Although we cannot directly observe general managers’ internal valuation models, the trade market aligns closely with what we would expect if teams were optimizing for elite player acquisition rather than expected performance value.
@@ 6.2 Surplus value
In Figure 15 we visualize expected surplus value (black dashed line) and the probability that surplus value exceeds a threshold (colored lines). The loser’s curse––top draft picks yield lower expected surplus value than those selected later in the first round––manifests as a spike in the expected surplus value curve above one during the first round. According to the probability that surplus value exceeds a high threshold, however, the effect size of the first-round spike diminishes and the surplus peak shifts earlier in the draft. Further, the posterior credible intervals are wide and cover one, suggesting these first-round spikes aren’t even significant.
@@ 6.3 Adjusting for position
In Figure 17 we visualize position-specific surplus value curves. For both quarterbacks and non-quarterbacks, expected surplus value peaks in the first round. According to the probability that surplus value exceeds a high threshold, however, there is no first-round peak for quarterbacks. Again, the posterior credible intervals are wide and cover one, suggesting these first-round spikes aren’t even significant. This suggests that, for teams seeking elite quarterbacks, the extreme upside of top selections offsets their higher cost. In contrast, even for elite non-quarterbacks, the first-round spike persists. Surplus value for non-quarterbacks peaks in the early-to-mid first round and declines steadily throughout the draft. The right tails of top-pick non-quarterbacks are not sufficiently fat to justify the high cost of early selections, suggesting that those picks are primarily priced for the ultra-high mean and variance of quarterbacks.
@@ 7 Discussion
A central premise in the judgment and decision making literature is that experts, even in high-stakes environments, exhibit systematic cognitive biases. Massey and Thaler, (2013) applied this framework to the NFL draft, concluding that, according to expected surplus value, general managers overvalue top draft picks, a finding widely cited as evidence of irrational decision making. This conclusion, however, hinges on the assumption that general managers should optimize for expected surplus value, an assumption that is neither explicitly justified nor necessarily aligned with the strategic realities of NFL team-building.
In this paper, we challenge that assumption and present an alternative framework for evaluating the value of a draft position. By estimating the full conditional density of player performance given draft position and player position, we construct draft position value curves from alternative objective functions. We focus on right tail probability––equivalently, the number of elite players expected to arise from a given pick––reflecting that general managers may be prioritizing the acquisition of elite, game-changing players. Under this perspective, top draft picks are not inefficiently overvalued but rather priced for their potential to produce transformational talent—players who can dramatically elevate a franchise’s trajectory. When valuing draft position based on right tail probability rather than expected surplus value, the first-round spike in the value of a draft position largely disappears. Teams selecting at the top of the draft retain higher relative value when their objective is to secure elite talent rather than merely optimizing expected surplus value. This effect is particularly pronounced for quarterbacks, whose performance distributions exhibit extremely fat right tails.

### v4

@@ 4 Research hypothesis
Massey and Thaler, (2013) developed the traditional analytical approach to constructing draft position value curves. Their study produced two key findings. First, they claimed to have found a loser’s curse––last years worst teams are “cursed” with the ostensible gifts of the top picks, which cost a lot (in compensation) without delivering enough on-field value to compensate. Second, they asserted that the draft pick trade market is highly inefficient––if one adopts the expected surplus value utility function, teams significantly overvalue top draft picks, consistently overpaying to trade up in the draft without trading down often enough.
Despite the widespread dissemination of these findings, general managers’ draft behavior has remained largely unchanged. When applied to recent data (2013––2023), the loser’s curse persists and the fitted trade market curve remains similar (see Section 3). If general managers were optimizing for expected performance value or expected surplus value, their continued reluctance to trade down and their willingness to pay a premium for higher picks would suggest systematic misjudgment. One possible explanation is that general managers have simply failed to learn from these analyses. An alternative explanation is that general managers are instead operating under a fundamentally different utility function. If maximizing expected surplus value does not align with their actual objectives, then their observed behavior may be rational, even if it appears suboptimal under Massey and Thaler’s framework. This paper challenges their core assumption, that general managers should adopt expected surplus value as their utility function in valuing draft picks. This assumption is neither explicitly justified in their analysis nor necessarily optimal in the context of NFL decision making.
@@ 5.1 Position-agnostic conditional density
In Figure 4 we visualize the empirical conditional density of performance $Y$ given draft position $x$. The empirical densities are extremely wobbly and noisy due to limited data for each individual draft position. But there is a general trend: we see a spike near zero that grows upwards as draft position increases and a right tail that shifts leftwards and diminishes as draft position increases. This motivates our “Bayesian spike plus Beta regression” model of ${\mathbb{P}}(Y|x)$.101010 We initially tried a lone Beta regression model without the spike near zero, which less successfully fit the right tail.
@@
We model the spike near zero, the bust spike, using logistic regression. We let a bust be a player whose performance outcome $Y$ is smaller than some cutoff ${y_{\text{bust}}}$. We model the bust probability ${\mathrm{bp}}(x):={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$ by
@@
where $\widetilde{x}$ is a B-spline basis.111111 The ${\mathsf{R}}$ code we used to generate the spline basis is splines::bs($x$, degree=3, df=4, intercept=T). We model the conditional precision by
@@
We use ${y_{\text{bust}}}=0.5\%$ and model the log-odds of the mean $\mu(x)$ using a spline, the log of the precision $\phi(x)$ with a line, and the log-odds of the bust probability ${\mathrm{bp}}(x)$ with a line because these parametrizations fit the data well––for instance, Figure 5 shows that the fitted curves closely track the trends of the empirical curves. Other values of ${y_{\text{bust}}}$ produced poorer fits (e.g., the fitted curves in Figure 5 tracked the empirical curves less closely). Modeling the precision with a spline led to overfitting, and using a spline for the bust probability produced negligible improvements over the linear model.
To capture uncertainty in the parameters, we use a Bayesian model, placing priors on the parameters. We put a weakly informative Normal prior with mean zero and standard deviation ten on each of the $\alpha$, $\beta$, and $\gamma$ parameters. Finally, we assume that $Y$ is uniformly distributed within the bust spike in a way that doesn’t vary with $x$: $Y|x,Y\leq{y_{\text{bust}}}\sim{\mathrm{Unif}}[0,{y_{\text{bust}}}]$.121212 We assess the reasonableness of this assumption later in this section.
@@
Figure 5: Left: our estimate of the conditional mean $x\mapsto\mu(x)={\mathbb{E}}[Y|x,Y>{y_{\text{bust}}}]$ given not a bust. Middle: our estimate of the conditional standard deviation $x\mapsto{\mathrm{sd}}(x)={\mathrm{sd}}(Y|x,Y>{y_{\text{bust}}})$ given not a bust. Right: our estimate of bust probability $x\mapsto{\mathrm{bp}}(x)={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$. The solid black lines are the posterior means, the cyan shaded regions are the $95\%$ credible intervals, and the gray dots are the empirical mean, empirical standard deviation, and empirical proportion values, respectively, for each draft position $x$.
Figure 6: The distribution of performance $Y$ ($x$-axis) given draft position $x$ (facet). The solid black lines are the fitted conditional densities ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})$ (the posterior mean curves) and the cyan shaded regions are the $95\%$ credible intervals, which overlay the empirical distributions (histograms and smoothed densities) in gray.
@@ 6 Results
Given our estimated conditional density of performance given draft position and player position, in this section we construct draft position value curves from alternative utility functions. We move beyond expected performance value and expected surplus value to utility functions that emphasize elite, right-tail outcomes.
@@ 6.1 Right tail probability
Further, we find that general managers trade draft picks as if they value a draft position by the probability it produces a player whose second contract exceeds $19.7\%$ of the salary cap. Specifically, the mean absolute error between the fitted trade market curve and a right tail probability curve with eliteness threshold $r$ is minimized by $r=19.7\%$ (note that Deshaun Watson has $r=19.7\%$). This is a high bar of eliteness: among players in our dataset who secured multi-year second contracts as regular starters, just $12.3\%$ of quarterbacks ($8/65$) and zero non-quarterbacks met this criterion. Although we cannot directly observe general managers’ internal valuation models, the trade market aligns closely with what we would expect if teams were optimizing for elite player acquisition rather than expected performance value.
@@ 6.2 Surplus value
In Figure 15 we visualize expected surplus value (black dashed line) and the probability that surplus value exceeds a threshold (colored lines). Massey and Thaler’s loser’s curse––top draft picks yield lower expected surplus value than those selected later in the first round––manifests as a spike in the expected surplus value curve above one during the first round.
Although the point estimate of the probability that surplus value exceeds a high threshold has a spike––though one that is smaller than and earlier in the draft than that of expected surplus value––consideration of the entire posterior distribution shows that a monotonically decreasing relationship is highly compatible with the data and model.
@@ 6.3 Adjusting for position
In Figure 17 we visualize position-specific surplus value curves. For both quarterbacks and non-quarterbacks, expected surplus value peaks in the first round.
Although the point estimate of the probability that surplus value exceeds a high threshold has a spike for non-quarterbacks, there is no spike for quarterbacks. Further, consideration of the entire posterior distribution for both quarterbacks and non-quarterbacks shows that a monotonically decreasing relationship is highly compatible with the data and model. Thus, particularly for teams seeking elite quarterbacks, our analysis provides strong support for the notion that the extreme upside of top selections offsets their higher cost.
@@ 7 Discussion
A central premise in the judgment and decision making literature is that experts, even in high-stakes environments, exhibit systematic cognitive biases. Massey and Thaler, (2013) applied this framework to the NFL draft, asserting that, according to expected surplus value, general managers overvalue top draft picks, a finding widely cited as evidence of irrational decision making. This assertion, however, hinges on the assumption that general managers should optimize for expected surplus value, an assumption that is neither explicitly justified nor necessarily aligned with the strategic realities of NFL team-building.
In this paper, we challenge that assumption and present an alternative framework for evaluating the value of a draft position. By estimating the full conditional density of player performance given draft position and player position, we construct draft position value curves from alternative utility functions. We focus on right tail probability––equivalently, the number of elite players expected to arise from a given pick––reflecting that general managers may be prioritizing the acquisition of elite, game-changing players. Under this perspective, top draft picks are not inefficiently overvalued but rather priced for their potential to produce transformational talent—players who can dramatically elevate a franchise’s trajectory. When valuing draft position based on right tail probability rather than expected surplus value, the first-round spike in the value of a draft position largely disappears. Teams selecting at the top of the draft retain higher relative value when their objective is to secure elite talent rather than merely optimizing expected surplus value. This effect is particularly pronounced for quarterbacks, whose performance distributions exhibit extremely fat right tails.